OpenStudy (anonymous):

If 3^(a+b) = 28 and 28^b = 81, find a in terms of b.

ganeshie8 (ganeshie8):

one way is to find the value of 28 from second equation :- 28^b = 81 28 = (81)^(1/b)

ganeshie8 (ganeshie8):

first equation says : 28 = 3^(a+b)

ganeshie8 (ganeshie8):

since left side is same, u can equate the right sides

ganeshie8 (ganeshie8):

(81)^(1/b) = 3^(a+b) see if u can massage this further :)

OpenStudy (loser66):

\(81=3^4\\3^{4/b}=3^{a+b}\\\dfrac{4}{b}=a+b\) isolate a .DAT SIT